OpenStudy (anonymous):

8x^2+7x-3=0 and 7c^2-2c=1.

OpenStudy (dumbcow):

use quadratic formula for solving ax^2 +bx+c = 0 \[x = \frac{-b \pm \sqrt{b^2 -4ac}}{2a}\]

OpenStudy (anonymous):

thank you!

OpenStudy (dumbcow):

yw :)

OpenStudy (anonymous):

i dont think im doing this right can you help me?

OpenStudy (anonymous):

@dumbcow

OpenStudy (dumbcow):

ok for #1 a = 8 b=7 c =-3 \[x = \frac{-7 \pm \sqrt{7^2 - 4(8)(-3)}}{2(8)}\]

OpenStudy (anonymous):

i got that part, but im getting weird numbers when i try and solve it

OpenStudy (dumbcow):

\[\frac{-7 \pm \sqrt{145}}{16}\] those weird numbers ? those are correct

OpenStudy (anonymous):

no i must have been doing it wrong. is that the answer?

OpenStudy (dumbcow):

yeah

OpenStudy (anonymous):

alright thanks so much

OpenStudy (dumbcow):

7^2 = 49 4*8*3 = 96 96+49 = 145

OpenStudy (anonymous):

can i ask you one more question?

OpenStudy (dumbcow):

sure

OpenStudy (anonymous):

nevermind i figured it out thanks

OpenStudy (dumbcow):

ok